OpenStudy (isabel☺):

solve y' = xe^(y-x^(2)), when x=0, y=0?

OpenStudy (isabel☺):

dy/dx right?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= xe^{y-x^2}\]

OpenStudy (anonymous):

\[xe^{y-x^2} = xe^y \times e^{-x^2}\]

OpenStudy (anonymous):

Might be able to separate the variables now and solve?

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = xe^y \times e^{-x^2}\] If we divide both sides by e^(y) and multiply both sides by (dx) \[\frac{ dy }{ e^y } = xe^{-x^2} dx\]

OpenStudy (anonymous):

Integrate both sides

OpenStudy (anonymous):

why integrate?

OpenStudy (anonymous):

To eliminate the dy and dx \[\int\limits_{}^{}\frac{ dy }{ e^y } = \int\limits_{}^{}xe^{-x^2} dx\] and so he can plug in the initial conditions to solve for C

OpenStudy (isabel☺):

ohh i thought that will be dy=x(e^y/ e^x^2)

OpenStudy (anonymous):

It is the same thing since there is a negative sign in the exponent of the e^-(x^2)

OpenStudy (isabel☺):

C=-1/2 ?

OpenStudy (anonymous):

Have not attempted the problem but if you show your work I can verify

OpenStudy (isabel☺):

\[\int\limits\limits \frac{ dy }{ e^y } - \int\limits\limits \frac{ xdx }{e ^{x ^{2}} } =0\] \[\int\limits \frac{ du }{ e^u } - \frac{ 1 }{ 2 } \int\limits \frac{ du }{ e ^{u} } = \int\limits 0\] \[\int\limits e ^{-u} du - \frac{ 1 }{ 2 } \int\limits e ^{-u} du = C\] \[\frac{ -1 }{ e ^{y} } + \frac{ 1}{ 2ex ^{2}} = C\] \[\frac{ 1 }{ 2 } e ^{-x ^{2}} = C + e ^{-y}\] \[\frac{ 1 }{ 2 } e ^{0} = C + e ^{0}\] \[1/2 = C+1 \] C=-1/2 ???

OpenStudy (isabel☺):

@Johnbc

OpenStudy (anonymous):

Great Job, that is correct. Using you're value of C plug it back into your final equation before you valued X and Y so that you have the full answer of y'(x).